OpenStudy (anonymous):

Write the equation of the line which passes through (2, 1) and is perpendicular to x = –2.

OpenStudy (anonymous):

for perpendicular condition. m=1/2 negative reciprocal of given value hence, y-1=(1/2)(x-2) y-1=(1/2)x-1 y=(1/2)x 2y=x

OpenStudy (anonymous):

if u have any query u can surely ask

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

u r welcome